OpenStudy (anonymous):

can some one help me please!!! The Smiths and Jones family each rented a ballroom for their wedding. The Smiths paid a $250 down payment plus $175 for each hour used. The Jones family paid a $400 down payment plus $125 per hour. If the weddings lasted the same amount of time and they paid the same amount at the end of the evening, how long were the weddings? a.2 b.3 c.4 b.5

OpenStudy (anonymous):

3 hours, I just had to use two calculators.

OpenStudy (anonymous):

how did you get that ? please

OpenStudy (anonymous):

Smiths: 175*3=525 525+250=775 --------------------------------- Jones: 125*3=375 375+400=775

OpenStudy (anonymous):

Now, do you get it. I tried to make it easier to understand, because it was sloppy before.

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Your Welcome :)